OpenStudy (anonymous):

For the function g(x) = 1 , find (g^-1 o g)(0). _____ x + 1

OpenStudy (anonymous):

g(x)=y=1/(x+1) x=1/y-(1) g^-1(x)=1/x-(1) g^-1o g(x)=1/(1/(1+x))-(1) = x g^-1 o g(o)=0